### v2

# Modular transformations of on-shell actions of (root-)$\overline{\text{T}}$ deformed holographic CFTs

Abstract
In this study, we examine the modular transformations of the (root-) $\overline{\text{T}}$ deformed torus partition function of a two-dimensional CFT (with a gravitational anomaly) from the holographic perspective by computing the on-shell actions of various saddle solutions of the dual gravity theories.
1State Key Laboratory of Quantum Optics and Quantum Optics Devices, Institute of Theoretical Physics, Shanxi University, Taiyuan 030006, P. R. China
2Kavli Institute for Theoretical Sciences (KITS),
University of Chinese Academy of Science, 100190 Beijing, P. R. China
3School of Physical Sciences, University of Chinese Academy of Sciences, Zhongguancun East Road 80, Beijing 100190, P. R. China
1 Introduction
In the past two decades, a very interesting irrelevant deformation of two-dimensional conformal field theories (CFTs), known as $\overline{\text{T}}$ deformation, was proposed [1, 2, 3]. One of the most intriguing properties of $\overline{\text{T}}$ deformed CFTs is that one can exactly calculate some quantities including energy spectrum [1, 2, 3] and S-Matrix [2, 4]. [^4] Moreover, the closed form of the Lagrangian of the $\overline{\text{T}}$ deformed theory can be obtained for some models, e.g. [5, 6]. In recent years, various features of the $\overline{\text{T}}$ deformed CFTs were investigated such as Lax connections[7, 8], S-Matrix [4, 9], correlation functions [11, 19, 10, 18, 12, 13, 14, 16, 20, 17, 26, 25, 15, 21, 22, 23, 24, 27], partition function[28, 29, 30, 32, 35, 33, 34, 31], entanglement entropy [36, 37, 38, 39, 41, 42, 40, 43, 55, 49, 50, 51, 52, 48, 46, 54, 53, 44, 45, 47, 56], pseudo entanglement entropy [57, 58], entanglement negativity [59], mutual information [61, 60, 52], entanglement wedge cross section [61, 60, 62, 56] and Action-Complexity [63, 65, 64]. Furthermore, theses deformations were extended to quantum field theories in zero dimension [66], one dimension [67, 68, 69], higher dimensions [71, 5, 70, 72, 73], to non-relativistic quantum field theories [74, 65, 6, 60] as well as to lattice theories [75]. See also [76, 77] which are very nice and comprehensive reviews on the recent developments.
On the other hand, there is another interesting deformation of CFTs dubbed Root-$\overline{\text{T}}$ deformation which is marginal and commutes with the $\overline{\text{T}}$ deformation [78, 79, 80, 82, 83, 81, 84, 85]. Moreover, similar to the $\overline{\text{T}}$ deformation, the deformation operator does not change under the flow [85]. However, it preserves the conformal invariance in the sense that the trace of the stress energy tensor remains zero under the flow [83, 85].
In the context of the AdS/CFT correspondence [86] , there are two different proposals for the holographic dual of $\overline{\text{T}}$ deformed CFTs: [^5] The holographic dual of root-$\overline{\text{T}}$ deformed CFTs was studied in [85]. cut-off proposal [87, 10] (or glue-on proposal [54, 88]) and mixed-boundary-condition proposal [89]. The cut-off proposal is easier to impliment, but it has several limitations[89]. On the other hand, the mixed-boundary-condition proposal is applicable to all holographic theories, regardless of the presence of matters, and the holographic dictionary can be summarized as follows: [77] 666the notation will become clear in a moment.
$Z_{\text{T$\overline{\text{T}}$ ,CFT}}[\gamma^{[\mu]}_{\alpha\beta}]=Z_{\text{grav}}\left[g_{\alpha\beta}^{(0)}+\frac{\mu}{16\pi G_{N}}g_{\alpha\beta}^{(2)}+\frac{\mu^{2}}{(16\pi G_{N})^{2}}g_{\alpha\beta}^{(4)}=\gamma^{[\mu]}_{\alpha\beta}\right],$ (1)
where $\mu$ is the deformation parameter. The key observation of us is that in the dictionary (1) we should add a proper boundary term to the gravity action [^7] Perhaps this is already hidden written in[89] and the presence of this them is stressed by one of the authors in [55].. The presence of this boundary term is natural considering that the double trace deformation not only shifts the source but also shifts the generating function [90, 91]. In this paper, we compute the on-shell actions of the $\overline{\text{T}}$ deformed theory defined on a torus. The aim is to derive and understand the modular properties of the $\overline{\text{T}}$ deformed torus partition function from the holographic perspective.
A two-dimensional CFT defined on a torus (and other Riemann surfaces) is subject to an important set of consistent conditions known as modular invariance. This requires that the partition function or correlation functions should remain the same, regardless of how the torus is represented. Modular invariance is a powerful constraint, leading to significant consequences such as the Hawking-Page transition [92] and the Cardy formula [93]. It has been discovered in [29] that the deformation parameter should also transform under the modular transformation which means that this kind of transformation can relate the partition functions of two different quantum field theories along the $\overline{\text{T}}$ flow. We find that this property can be easily understood by observing that the deformation parameter $\mu$ has the same dimension of the radial coordinate and undergoes rescaling under a Weyl transformation. Meanwhile, we derive the thermodynamic first law and Smarr relations for the deformed BTZ black hole. Furthermore, we show that the modular property remains preserved even in the CFT with a gravitational anomaly[94], where the holomorphic sector and anti-holomorphic sector have different central charges[95], which is dual to a topologically massive gravity (TMG) theory [96, 97]. As a contrast, we also consider a marginal deformation, the root-$\overline{\text{T}}$ deformation [82, 83] in the mixed-boundary-condition proposal [85], for which the deformation parameter is dimensionless and we find indeed it does not transform under modular transformation.
2 Mixed-boundary-condition proposal and the deformed on-shell action
In this section, we briefly review the mixed-boundary-condition proposal of the $\overline{\text{T}}$ deformation with a special emphasis on the deformed on-shell action. The $\overline{\text{T}}$ deformation of 2d field theory with action $S_{\text{CFT}}$ is defined by the flow equation
$\frac{dS^{[\mu]}_{\text{CFT}}}{d\mu}=\int d^{2}x\sqrt{\gamma}\,\text{T}\bar{\text{T}}^{[\mu]},$ (2.1)
where the double trace operator $\text{T}\bar{\text{T}}^{[\mu]}$ is defined as
$\text{T}\bar{\text{T}}^{[\mu]}=\frac{1}{8}\left(T_{\alpha\beta}T^{\alpha\beta}-(T^{\alpha}_{\alpha})^{2}\right),$ (2.2)
and it is evaluated in the deformed theory with deformation parameter $\mu$. In the linear order, the deformation is a special double trace deformation
$S_{CFT}^{[\mu]}=S_{CFT}+\mu\int d^{2}x\sqrt{\gamma}\,\text{T$\overline{\text{T}}$ }+\mathcal{O}(\mu^{2}).$ (2.3)
The specialty of this double trace deformation is that the operator $\overline{\text{T}}$ depends both on the boundary metric and the energy-stress tensor. Consequently, both of them flow under the deformation according to [89]:
$\gamma_{\alpha\beta}^{[\mu]}=\gamma_{\alpha\beta}^{[0]}+\frac{1}{2}\mu\hat{T}_{\alpha\beta}^{[0]}+\frac{1}{16}\mu^{2}\hat{T}_{\alpha\rho}^{[0]}\hat{T}_{\sigma\beta}^{[0]}{\gamma^{[0]}}^{\rho\sigma},$ (2.4)
$\hat{T}^{[\mu]}_{\alpha\beta}=\hat{T}^{[0]}_{\alpha\beta}+\frac{1}{4}\mu\hat{T}_{\alpha\rho}^{[0]}\hat{T}_{\sigma\beta}^{[0]}{\gamma^{[0]}}^{\rho\sigma},$ (2.5)
where $\hat{T}_{\alpha\beta}=T_{\alpha\beta}-\gamma_{\alpha\beta}T$, see Appendix A for details. The initial data $(\gamma_{\alpha\beta}^{[0]},\hat{T}^{[0]}_{\alpha\beta})$ defines the seed CFT which generally should not be viewed as the undeformed CFT because from the field theory point of view the $\overline{\text{T}}$ deformation only deforms the theory but not the spacetime background [^8] Do not be confused with the dynamical coordinate transformation interpretation [98, 99] of the $\overline{\text{T}}$ deformation.. The seed CFT is dual to an asymptotic AdS3 solution which can be written in the Fefferman-Graham gauge [100]
$ds^{2}=g_{\alpha\beta}(z,x^{\alpha})dx^{\alpha}dx^{\beta}+\frac{dz^{2}}{z^{2}},\quad g_{\alpha\beta}(z,x^{\alpha})=\frac{g_{\alpha\beta}^{(0)}}{z^{2}}+g_{\alpha\beta}^{(2)}+z^{2}g_{\alpha\beta}^{(4)},\quad g_{\alpha\beta}^{(0)}=\gamma_{\alpha\beta}^{[0]},$ (2.6)
where $g^{(2)}$ corresponds to the expectation value of the seed CFT operator
$\hat{T}^{[0]}_{\alpha\beta}=\frac{1}{8\pi G_{N}}g_{\alpha\beta}^{(2)}.$ (2.7)
The solution (2.4) implies that the deformed boundary metric can be interpreted as the induced metric at the cut-off surface $z=z_{c}\equiv\sqrt{\rho_{c}}$:
$\gamma_{\alpha\beta}^{[\mu]}=g_{\alpha\beta}^{(0)}+\frac{\mu}{16\pi G_{N}}g_{\alpha\beta}^{(2)}+\frac{\mu^{2}}{(16\pi G_{N})^{2}}g_{\alpha\beta}^{(4)}=\rho_{c}g_{\alpha\beta}(z_{c}=\sqrt{\rho_{c}}),\quad\rho_{c}=\frac{\mu}{16\pi G_{N}}.$ (2.8)
We will interchangeably use $\rho_{c}$ and $\mu$ as the deformation parameters.
As pointed out in [55], the proper on-shell action of the $\overline{\text{T}}$ deformed theory should be
$I_{\text{on-shell}}^{[\mu]}$ $=$ $I_{\text{Euclidean}}^{[0]}\left(\gamma^{[\mu]}_{\alpha\beta}=g_{\alpha\beta}^{(0)}+\frac{\mu}{16\pi G_{N}}g_{\alpha\beta}^{(2)}+\frac{\mu^{2}}{(16\pi G_{N})^{2}}g_{\alpha\beta}^{(4)}\right)-\mu\int\sqrt{\gamma^{[\mu]}}\text{T}\bar{\text{T}}^{[\mu]}$ (2.9)
$\equiv$ $I_{\text{bulk}}^{[\mu]}+I_{\text{bdy}}^{[\mu]}.$
Using (2.4) and (2.5) one can show that $\sqrt{\gamma}\text{T$\overline{\text{T}}$ }$ is invariant under the flow so the boundary term in (2.9) can also be written as
$-\mu\int\sqrt{\gamma^{[\mu]}}\text{T}\bar{\text{T}}^{[\mu]}=-\mu\int\sqrt{g^{[0]}}\text{T}\bar{\text{T}}^{[0]}.$ (2.10)
When the seed metric $g_{\alpha\beta}^{(0)}$ is flat, the bulk solution in the FG gauge is characterized by two arbitrary functions as
$ds^{2}=\frac{dz^{2}}{z^{2}}+\frac{1}{z^{2}}(dv+\bar{L}_{0}(\bar{v})z^{2}d\bar{v})(d\bar{v}+L_{0}(v)z^{2}dv).$ (2.11)
In this case, one can show that the deformed metric $\gamma_{\alpha\beta}^{[\mu]}$ is also flat thus (2.8) is related to $g_{\alpha\beta}^{(0)}$ via the coordinate transformations
$dVd\bar{V}=\left(dv+\bar{L}_{0}(\bar{v})\rho_{c}d\bar{v}\right)\left(d\bar{v}+{L}_{0}({v})\rho_{c}d{v}\right),$ (2.12)
[…]
Substituting the inverse of (2.13) into (2.11) leads to the metric of the dual geometry of the $\overline{\text{T}}$ deformed CFT
$ds_{\mu}^{2}$ $=$ $\frac{dz^{2}}{z^{2}}+\frac{1}{z^{2}(1-\rho_{c}^{2}L_{0}\bar{L}_{0})^{2}}\left((1-\rho_{c}L_{0}\bar{L}_{0}z^{2})dV+\bar{L}_{0}(z^{2}-\rho_{c})d\bar{V}\right)$
$\left((1-\rho_{c}L_{0}\bar{L}_{0}z^{2})d\bar{V}+{L}_{0}(z^{2}-\rho_{c})d{V}\right),$ (2.14)
where $L_{0}$ and $\bar{L}_{0}$ should be understood as functions of $V$ and $\bar{V}$. The deformed stress tensor is given by (2.5):
$T_{\alpha\beta}^{[\mu]}dx^{\alpha}dx^{\beta}=\frac{L_{0}dV^{2}+\bar{L}_{0}d\bar{V}^{2}+2\rho_{c}L_{0}\bar{L}_{0}dVd\bar{V}}{8\pi G_{N}(1-\rho_{c}^{2}L_{0}\bar{L}_{0})},$ (2.15)
and the deformed on-shell action is equal to [55]
$I_{\text{on-shell}}^{[\mu]}=-\frac{c}{6\pi}\int\sqrt{g^{[0]}}\left(\sqrt{L_{0}\bar{L}_{0}}+\rho_{c}L_{0}\bar{L}_{0}\right),$ (2.16)
where $c=3/(2G_{N})$ [101]. Since we are interested in the modular property of the torus partition function, below we will assume that the boundary manifold is a flat torus with two generic complex periods $\bm{\omega}=\omega_{1}+\text{i}\omega_{2}$ and $\bm{\beta}=\beta_{1}+\text{i}\beta_{2}$ so the possible bulk solutions are global AdS, BTZ black hole and other $SL(2,R)$ black holes. Because AdS3 gravity does not have local degrees of freedom, these solutions are all diffeomorphic equivalent but distinguished by the periodicity. A torus is specified by the modular parameter defined by the ratio of the two complex periods: $\bm{\tau}=\frac{\bm{\beta}}{\bm{\omega}}$. It is well known that the undeformed on-shell actions of these solutions are related to each other via the $SL(2,\mathbb{Z})$ modular transformation. This modular property can be traced back to the diffeomorphic equivalence of bulk solutions. However, it may be broken due to the non-commutativity of the $\overline{\text{T}}$ deformation with the diffeomorphism. Nevertheless, [29] shows that the $\overline{\text{T}}$ deformed torus partition function exhibits a modified modular invariance:
$Z\left(\bm{\tau}\Big{|}\rho_{c}\right)=Z\left(\frac{a\bm{\tau}+b}{c\bm{\tau}+d}\Big{|}\frac{\rho_{c}}{|c\bm{\tau}+d|^{2}}\right),$ (2.17)
by directly calculating the torus partition function of a $\overline{\text{T}}$ deformed 2d CFT. In this work, we aim to understand this modular property holographically by computing the deformed on-shell actions of the dual gravity theory. Before turning on the $\overline{\text{T}}$ deformation, we will construct the global AdS and black hole solutions with two complex periods $\bm{\omega}$ and $\bm{\beta}$ as a prerequisite.
We want to emphasize that during the $\overline{\text{T}}$ deformation process, three distinct CFTs are involved: the seed CFT defined on the seed torus with metric $ds^{2}=dvd\bar{v}$, the deformed CFT defined on the deformed torus with metric $ds^{2}=dVd\bar{V}$, and the undeformed CFT defined on the undeformed torus with metric $d\tilde{s}^{2}=d\tilde{v}d\bar{\tilde{v}}$. The explicit relationships between these three CFTs will be discussed in each of the subsequent examples.
3 General global AdS
Let us start from the standard global AdS solution with [^9] In this work, we set the AdS radius to be $\ell=1$.
$L_{0}=\bar{L}_{0}=-\frac{1}{4},\quad\bm{\omega}=2\pi,\quad\bm{\beta}=2\pi\bm{\tau},$ (3.1)
whose metric is then
$ds^{2}=\frac{dz^{2}}{z^{2}}+\frac{(z^{2}-4)^{2}d\phi^{2}}{16z^{2}}+\frac{(z^{2}+4)^{2}dt^{2}}{16z^{2}},\quad\phi+\text{i}t\equiv v\,,$ (3.2)
which can be transformed into a more familiar form by introducing the radial coordinate $r=|\frac{(z^{2}-4)}{4z}|.$ The modular parameter $\bm{\tau}$ can be chosen freely and later on we will use this freedom to construct other $SL(2,\mathbb{Z})$ black holes following [102]. A global AdS with a general $\bm{\omega}$ can be obtained by first rescaling the coordinate $v$ as
$v\rightarrow\frac{2\pi}{\bm{\omega}}v,$ (3.3)
and then adjusting the radial coordinate $z$ accordingly such that the metric is in the standard form (2.11). We find that
${L}_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}},\quad\bar{{L}}_{0}=-\frac{\pi^{2}}{\bar{\bm{\omega}}^{2}}.$ (3.4)
and the resulting metric is
$ds^{2}=\frac{dz^{2}}{{z}^{2}}+g_{{t}{t}}d{t}^{2}+\frac{(1+{L}_{0}{z}^{2})(1+\bar{{L}}_{0}{z}^{2})}{{z}^{2}}(d{\phi}+\Omega d{t})^{2},$ (3.5)
where
$g_{{t}{t}}=\frac{(1-{z}^{4}{L}_{0}\bar{{L}}_{0})^{2}}{{z}^{2}(1+{L}_{0}{z}^{2})(1+\bar{{L}}_{0}{z}^{2})},\quad\Omega=\frac{\text{i}(\bar{{L}}_{0}-{L}_{0}){z}^{2}}{(1+{L}_{0}{z}^{2})(1+\bar{{L}}_{0}{z}^{2})}.$ (3.6)
Note that the metric will have no horizons when ${L}_{0}=\bar{{L}}_{0}$ which means $\bm{\omega}=\bar{\bm{\omega}}$.
3.1 $\overline{\text{T}}$ deformation
According to the dictionary (2.13), the $\overline{\text{T}}$ deformed theory is also defined on a torus with a flat metric $ds^{2}=dVd\bar{V}$, which is related to seed metric via
$V=v+\bar{L}_{0}\rho_{c}\bar{v},\quad\bar{V}=\bar{v}+L_{0}\rho_{c}v,$ (3.7)
$v=\frac{V-\rho_{c}\bar{L}_{0}\bar{V}}{1-\rho_{c}^{2}L_{0}\bar{L}_{0}},\quad\bar{v}=\frac{\bar{V}-\rho_{c}{L}_{0}{V}}{1-\rho_{c}^{2}L_{0}\bar{L}_{0}},$ (3.8)
where $L_{0}$ and $\bar{L}_{0}$ are given by (3.4). It implies that the two periods of the deformed torus are
$\bm{\omega}_{\mu}=\bm{\omega}+\bar{L}_{0}\rho_{c}{\bm{\bar{\omega}}}=\bm{\omega}(1-\frac{\pi^{2}\rho_{c}}{\bm{\omega}^{2}}),\quad\bm{\beta}_{\mu}=\bm{\beta}-\frac{\pi^{2}\rho_{c}}{\bm{\omega}^{2}}\bm{\bar{\beta}},$ (3.9)
and the $\overline{\text{T}}$ deformed bulk geometry is described by the metric
$ds_{\mu}^{2}=\frac{dz^{2}}{z^{2}}-\frac{\pi^{2}}{\bm{\omega}^{2}}(dv^{2}+d\bar{v}^{2})+\frac{(\bm{\omega}^{4}+\pi^{4}z^{4})dvd\bar{v}}{\bm{\omega}^{4}z^{2}},$ (3.10)
with (3.8). Using (2.15) we find that the deformed holographic mass and momentum are
$M_{\mu}$ $=$ $\frac{c\bm{\omega}_{\mu}}{6\pi}\frac{L_{0}}{1+\rho_{c}L_{0}}=-\frac{c\pi}{3(\bm{\omega}_{\mu}+\sqrt{\bm{\omega}_{\mu}^{2}+4\pi^{2}\rho_{c}})},$ (3.11)
$J_{\mu}$ $=$ $0,$ (3.12)
Defining $M_{0}=M_{\mu}|_{\mu=0}=-\frac{c\pi}{6\bm{\omega}_{\mu}}$, we find that the deformed energy can be rewritten as
$M_{\mu}=\frac{2\bm{\omega}_{\mu}}{\mu}\left(1-\sqrt{1-\mu\frac{{M}_{0}}{\bm{\omega}_{\mu}}}\right),\quad\mu={16\pi G_{N}}\rho_{c}\,,$ (3.13)
which is consistent with the field theory calculation of the deformed energy [1, 2, 3]. It may be confusing that the undeformed energy $M_{0}$ depends on $\bm{\omega}_{\mu}$ which depends on the deformation parameter $\rho_{c}$ implicitly via (3.9). However, as we emphasized in last section that when we compare the deformed CFT and the undeformed one we should fix the size of the system. That is why the undeformed energy $M_{0}$ can be obtained directly by setting $\rho_{c}=0$ in (3.11).
The deformed on-shell action can be easily worked out by substituting (3.4) into (2.16) and rewriting $(\bm{w},\bm{\beta})$ in terms of $(\bm{\omega}_{\mu},\bm{\beta}_{\mu}$):
$I_{E}^{[\mu]}$ $=$ $-\frac{c\pi A_{\mu}}{3}\frac{1}{|\bm{\omega}_{\mu}|^{2}+|\bm{\omega}_{\mu}|\sqrt{4\pi^{2}\rho_{c}+|\bm{\omega}_{\mu}|^{2}}}$
$=$ $-\frac{cA_{\mu}}{6\pi}\frac{\pi^{2}}{|\bm{\omega}_{\mu}|^{2}}{+}\rho_{c}\frac{cA_{\mu}}{6\pi}\frac{\pi^{4}}{|\bm{\omega}_{\mu}|^{4}}+\mathcal{O}(\rho_{c}^{2}),$
$=$ $I_{E}^{[0]}[\bm{\omega}_{\mu}]+\mu\int d^{2}x\sqrt{\gamma}\,\text{T$\overline{\text{T}}$ }+\mathcal{O}(\mu^{2}),$ (3.15)
where $A_{\mu}=|\bm{\omega}_{\mu}|^{2}\text{Im}(\bm{\tau}_{\mu})$ is the area of the deformed torus. In particular, the on-shell action equals to the free energy
$I_{E}^{[\mu]}=\text{Im}(\bm{\beta}_{\mu})M_{\mu},$ (3.16)
and the expansion (3.15) coincides exactly with the field theory result (2.3).
4 The BTZ black hole
An Euclidean rotating BTZ black hole has the standard metric
$ds^{2}=\frac{(r^{2}-R_{+}^{2})(r^{2}+R_{-}^{2})}{r^{2}}dt^{2}+\frac{r^{2}}{(r^{2}-R_{+}^{2})(r^{2}+R_{-}^{2})}dr^{2}+r^{2}(d\phi-\frac{R_{+}R_{-}}{r^{2}}dt)^{2},$ (4.1)
where its two (complex) periods
$(\phi,t)\sim(\phi+2\pi,t),\quad(\phi,t)\sim(\phi+\theta,t+\beta),$ (4.2)
are related to the inner and outer horizon radii via
$\beta=\frac{2\pi R_{+}}{R_{+}^{2}+R_{-}^{2}},\quad\theta=\frac{2\pi R_{-}}{R_{+}^{2}+R_{-}^{2}}.$ (4.3)
To obtain a BTZ black hole solution with two generic complex periods $(\bm{\omega}=\omega_{1}+\text{i}\omega_{2},\bm{\beta}=\beta_{1}+\text{i}\beta_{2})$, we can perform a similar rescaling as (3.3) to the standard BTZ metric (4.1). Transforming the resulting metric into the FG gauge, we find that
$L_{0}=-\frac{\pi^{2}}{\bm{\beta}^{2}},\quad\bar{L}_{0}=-\frac{\pi^{2}}{\bm{\bar{\beta}}^{2}}.$ (4.4)
For more details refer to Appendix B. Because the effect of the rescaling is like a boost (or more precisely a rotation in the Euclidean signature), the energy and the momentum get transformed accordingly and take a slightly complicated form
$M$ $=$ $-\frac{c\pi}{6|\bm{\beta}|^{4}}\left(\omega_{1}(\beta_{1}^{2}-\beta_{2}^{2})+2\omega_{2}\beta_{1}\beta_{2}\right),$ (4.5)
$J$ $=$ $-\frac{c\pi}{6|\bm{\beta}|^{4}}\left(\omega_{2}(\beta_{1}^{2}-\beta_{2}^{2})-2\omega_{1}\beta_{1}\beta_{2}\right),$ (4.6)
while the black hole entropy is invariant under the boost and manifestly it only depends on the modular parameter
$S_{BH}=\frac{c\pi}{3}\frac{\text{Im}(\bm{\tau})}{|\bm{\tau}|^{2}}.$ (4.7)
The details of the these deviations can be found in Appendix B. It is straightforward to check that the Smarr relation
$S_{BH}=2(\beta_{2}M+\beta_{1}J),$ (4.8)
and the Cardy formula
$S_{BH}=\sqrt{\frac{c\pi\bm{\omega}}{6}{(M+\text{i}J)}{}}+\sqrt{\frac{c\pi\bm{\bar{\omega}}}{6}{(M-\text{i}J)}{}}$ (4.9)
are satisfied.
4.1 The $\overline{\text{T}}$ deformation
The $\overline{\text{T}}$ deformed theory is also defined on a torus with metric $ds^{2}=dVd\bar{V}$ where the deformed coordinates $(V,\bar{V})$ are related to seed metric via (3.7) and (3.8), however now $L_{0}$ and $\bar{L}_{0}$ are given by (4.4). It implies that the two complex periods of the deformed torus are
$\bm{\omega}_{\mu}=\bm{\omega}-\frac{\pi^{2}\rho_{c}}{\bm{\bar{\beta}}^{2}}\bm{\bar{\omega}},\quad\bm{\beta}_{\mu}=\bm{\beta}(1-\frac{\pi^{2}\rho_{c}}{|\bm{\beta}|^{2}}),$ (4.10)
and the $\overline{\text{T}}$ deformed bulk geometry is described by the corresponding metric (2). The relation (4.10) implies that $\bm{\omega}_{\mu}$ is complex in general and the geometry describes a general deformed $SL(2,\mathbb{Z})$ black hole which we will consider later on. For BTZ solutions, we choose $\bm{\omega}_{\mu}$ to be real. The deformed energy and momentum can be found directly by using (2.15).
The energy and the momentum of the deformed theory are given by
$M_{\mu}$ $=$ $\frac{c\bm{\omega}_{\mu}}{12\pi}\frac{1}{\rho_{c}}\left(1-\frac{|\bm{\beta}_{\mu}|}{\sqrt{|\bm{\beta}_{\mu}|^{2}+4\pi^{2}\rho_{c}}}\frac{|\bm{\beta}_{\mu}|^{4}+(\bm{\beta}_{\mu}+\bm{\bar{\beta}}_{\mu})^{2}\pi^{2}\rho_{c}}{|\bm{\beta}_{\mu}|^{4}}\right),$ (4.11)
$J_{\mu}$ $=$ $\frac{c\pi\bm{\omega}_{\mu}}{12}\frac{\text{i}(\bm{\bar{\beta}}_{\mu}^{2}-\bm{\beta}_{\mu}^{2})}{|\bm{\beta}_{\mu}|^{3}\sqrt{|\bm{\beta}_{\mu}|^{2}+4\pi^{2}\rho_{c}}}.$ (4.12)
In the limit $\rho_{c}=0$, they reduce to
$M_{0}$ $=$ $M_{\mu}\Big{|}_{\mu=0}=-\frac{c\pi\bm{\omega}_{\mu}}{12}\left(\frac{1}{\bm{\beta}_{\mu}^{2}}+\frac{1}{\bm{\bar{\beta}}_{\mu}^{2}}\right),$ (4.13)
$J_{0}$ $=$ $J_{\mu}\Big{|}_{\mu=0}=-\frac{c\pi\bm{\omega}_{\mu}}{12}\text{i}\left(\frac{1}{\bm{\bar{\beta}}_{\mu}^{2}}-\frac{1}{\bm{\beta}_{\mu}^{2}}\right).$ (4.14)
However, $M_{0}$ and $J_{0}$ are not the undeformed energy and momentum. To see this, for example by setting $J_{0}=J_{\mu}=0$ (i.e. $\bm{\beta}_{\mu}=-\bm{\bar{\beta}}_{\mu}$) we find
$M_{\mu}=\frac{2\bm{\omega}_{\mu}}{\mu}\left(1-\frac{1}{\sqrt{1+\mu\frac{M_{0}}{\bm{\omega}_{\mu}}}}\right),$ (4.15)
which is different from the field theory expectation (3.13). The mismatch is due to the fact that the $\overline{\text{T}}$ deformation alters the temperature of the system! In fact, the undeformed CFT lives on the torus with the flat metric $d\tilde{s}^{2}=d\tilde{v}d\bar{\tilde{v}}$ and coordinate $\tilde{v}$ is defined by [89]
$\tilde{v}=\frac{V-\rho_{c}\bar{L}_{0}\bar{V}}{1-\rho_{c}\bar{L}_{0}}=\frac{1-\rho_{c}^{2}L_{0}\bar{L}_{0}}{1-\rho_{c}\bar{L}_{0}}v,\quad\bar{\tilde{v}}=\frac{\bar{V}-\rho_{c}{L}_{0}{V}}{1-\rho_{c}{L}_{0}}=\frac{1-\rho_{c}^{2}L_{0}\bar{L}_{0}}{1-\rho_{c}{L}_{0}}\bar{v},$ (4.16)
where $L_{0}$ and $\bar{L}_{0}$ are given by (4.4). Therefore, the two complex periods of this undeformed torus are
$\bm{\tilde{\omega}}=\bm{\omega}_{\mu},\quad\bm{\tilde{\beta}}=\bm{\beta}_{\mu}\frac{1+\frac{\rho_{c}\pi^{2}}{|{\bm{\beta}}|^{2}}}{1+\frac{\rho_{c}\pi^{2}}{\bm{\bar{\beta}}^{2}}}=\frac{\left|\bm{\beta}_{\mu}\right|^{2}\sqrt{\left|\bm{\beta}_{\mu}\right|^{2}+4\pi^{2}\rho_{c}}}{\text{Re}\left(\bm{\beta}_{\mu}\right)\sqrt{\left|\bm{\beta}_{\mu}\right|^{2}+4\pi^{2}\rho_{c}}-\text{i}\left|\bm{\beta}_{\mu}\right|\text{Im}\left(\bm{\beta}_{\mu}\right)}.$ (4.17)
When $\bm{\beta}_{\mu}=\text{i}\beta_{\mu}$ is purely imaginary, so are $\bm{\beta}=\text{i}\beta$ and $\bm{\tilde{\beta}}=\text{i}\tilde{\beta}$ then the relation (4.17) is simplified to
$\tilde{\beta}=\sqrt{\beta_{\mu}^{2}+4\pi^{2}\rho_{c}}.$ (4.18)
This coordinate transformation (4.16) is uniquely fixed by requiring that it does not change the horizon area (black hole entropy) and the momentum which are supposed to be invariant under $\overline{\text{T}}$ flow. Clearly, the undeformed CFT and the deformed CFT have the same spatial dimension. Note that the undeformed CFT and the seed CFT is related by a Weyl transformation:
$dvd\bar{v}=\frac{(1-\rho_{c}L_{0})(1-\rho_{c}\bar{L}_{0})}{(1-\rho_{c}^{2}L_{0}\bar{L}_{0})^{2}}d\tilde{v}d\bar{\tilde{v}}\equiv W^{2}d\tilde{v}d\bar{\tilde{v}},\quad W=\frac{|1+\frac{\rho_{c}\pi^{2}}{\bm{\beta}^{2}}|}{1-\frac{\rho_{c}^{2}\pi^{4}}{|\bm{\beta}|^{4}}},$ (4.19)
therefore, the undeformed energy and momentum are
$\tilde{M}$ $=$ $\frac{c\tilde{\bm{\omega}}}{12\pi}\frac{\left(L_{0}+\bar{L}_{0}\right)\left(1+L_{0}\bar{L}_{0}\rho_{c}^{2}\right)-4\rho_{c}L_{0}\bar{L}_{0}}{\left(1-L_{0}\bar{L}_{0}\rho_{c}^{2}\right)},$ (4.20)
$=$ $-\frac{c\pi\bm{\omega}_{\mu}}{3}\frac{(\bm{\beta}_{\mu}+\bm{\bar{\beta}}_{\mu})^{2}(\left|\bm{\beta}_{\mu}\right|^{2}+2\pi^{2}\rho_{c})-2\left|\bm{\beta}_{\mu}\right|^{4}}{\left|\bm{\beta}_{\mu}\right|^{3}\sqrt{\left|\bm{\beta}_{\mu}\right|^{2}+4\pi^{2}\rho_{c}}\left(\left|\bm{\beta}_{\mu}\right|+\sqrt{\left|\bm{\beta}_{\mu}\right|^{2}+4\pi^{2}\rho_{c}}\right)^{2}},$
$\tilde{J}$ $=$ $J_{\mu},$ (4.21)
and they satisfy
$M_{\mu}=\frac{2\tilde{\bm{\omega}}}{\mu}\left[1-\sqrt{1-\frac{\mu\tilde{M}}{\tilde{\bm{\omega}}}-\frac{\mu^{2}\tilde{J}^{2}}{4\tilde{\bm{\omega}}^{2}}}\;\right],$ (4.22)
as expected from the field theory result. Substituting (4.4) into (2.16) and using the relation (4.10), we find that the deformed on-shell action of the BTZ black hole is
$I_{E}^{[\mu]}$ $=$ $-\frac{c\pi A_{\mu}}{3}\frac{1}{|\bm{\beta}\mu|^{2}+|\bm{\beta}_{\mu}|\sqrt{4\pi^{2}\rho_{c}+|\bm{\beta}_{\mu}|^{2}}}$
$=$ $-\frac{cA_{\mu}}{6\pi}\frac{\pi^{2}}{|\bm{\beta}_{\mu}|^{2}}{+}\rho_{c}\frac{cA_{\mu}}{6\pi}\frac{\pi^{4}}{|\bm{\beta}_{\mu}|^{2}}+\mathcal{O}(\rho_{c}^{2}),$
$=$ $I_{E}^{(0)}+\mu\int d^{2}x\sqrt{\gamma}\,\text{T$\overline{\text{T}}$ }+\mathcal{O}(\mu^{2}),$ (4.24)
and the expansion (4.24) coincides exactly with the field theory result (2.3), where $A_{\mu}=|\bm{\omega}_{\mu}|^{2}\text{Im}(\bm{\tau}_{\mu})$ is the area of the deformed torus.
4.2 Thermodynamics
After the $\overline{\text{T}}$ deformation, the entropy of the black hole does not get deformed, thus it is equal to101010which is also equal to the black hole entropy of the undeformed theory.
$S_{BH}^{[\mu]}=\frac{c\pi}{3}\frac{A_{0}}{|\bm{\beta}|^{2}}=\frac{c\pi}{3}\frac{A_{\mu}}{|\bm{\beta}_{\mu}|^{2}}\frac{1}{\sqrt{1+\frac{4\pi^{2}\rho_{c}}{|\bm{\beta}_{\mu}|^{2}}}}=\frac{c\pi}{3}\frac{\text{Im}(\bm{\tau}_{\mu})}{|\bm{\tau}_{\mu}|^{2}\sqrt{1+\frac{4\pi^{2}}{|\bm{\tau}_{\mu}|^{2}}\frac{\rho_{c}}{|\bm{\omega}_{\mu}|^{2}}}},$ (4.25)
which not only depends on the modular parameter but also depends on the size of the system. For example, the deformation is insignificant when the torus is infinite long. This is because that the effective deformation parameter is the dimensionless quantity $\rho_{c}/|\bm{\omega}_{\mu}|^{2}$. As expected, the deformed on-shell action still equals to the free energy
$I_{E}^{[\mu]}=\text{Im}(\bm{\beta}_{\mu})M_{\mu}+\text{Re}(\bm{\beta}_{\mu})J_{\mu}-S_{BH}^{[\mu]},$ (4.26)
even though both sides of this equation are deformed in a non-trivial way. To examine the first law of thermodynamics we treat $M_{\mu}$, $J_{\mu}$ and $S_{BH}^{[\mu]}$ as functions of $\bm{\beta}_{\mu}$ and $\bm{\bar{\beta}}_{\mu}$ and do the variation with respect to them. It turns out the first law is in the same form of the undeformed one
$\text{Im}(\bm{\beta}_{\mu})\delta M_{\mu}+\text{Re}(\bm{\beta}_{\mu})\delta J_{\mu}=\delta S_{BH}^{[\mu]},$ (4.27)
but the Smarr relation takes a different form
$\text{Im}(\bm{\beta}_{\mu})M_{\mu}+\text{Re}(\bm{\beta}_{\mu})J_{\mu}=S_{BH}^{[\mu]}\frac{1}{1+\sqrt{1+\frac{4\pi^{2}\rho_{c}}{|\bm{\beta}_{\mu}|^{2}}}}.$ (4.28)
Interestingly, the Smarr relation looks like the standard one if we rewrite $\bm{\beta}_{\mu}$ in terms of $\bm{\beta}$ via (4.10)
$\text{Im}(\bm{\beta})M_{\mu}+\text{Re}(\bm{\beta})J_{\mu}=\frac{1}{2}S_{BH}^{[\mu]}.$ (4.29)
On the other hand, the Cardy formula is now
$S_{BH}^{[\mu]}=\sqrt{\frac{c\pi\bm{\omega}_{\mu}}{6}{(M_{\mu}+\text{i}J_{\mu})}{}-\frac{c\pi\mu}{24}(M_{\mu}^{2}+J_{\mu}^{2})}+\sqrt{\frac{c\pi\bm{\bar{\omega}_{\mu}}}{6}{(M_{\mu}-\text{i}J_{\mu})}{}-\frac{c\pi\mu}{24}(M_{\mu}^{2}+J_{\mu}^{2})}.$
[…]
$S_{BH}^{[\mu]}$ $=$ ${\lim_{n\rightarrow 1}}\frac{1}{1-n}\log\frac{Z_{n}}{Z_{1}^{n}}={\lim_{n\rightarrow 1}}\frac{1}{1-n}\log\frac{e^{-I^{[\mu]}_{\text{on-shell,BTZ}}(n\bm{\tau}_{\mu})}}{e^{-nI^{[\mu]}_{\text{on-shell,BTZ}}(\bm{\tau}_{\mu})}}$ (4.31)
$=$ ${\lim_{n\rightarrow 1}}\frac{cn\,\text{Im}(\bm{\tau}_{\mu})|\bm{\omega}_{\mu}^{2}|\left(\sqrt{1+\frac{4\pi^{2}\rho_{c}}{|\bm{\omega}_{\mu}\bm{\tau}_{\mu}|^{2}}}-\sqrt{1+\frac{4\pi^{2}\rho_{c}}{|\bm{\omega}_{\mu}\bm{\tau}_{\mu}|^{2}n^{2}}}\right)}{12(n-1)\pi\rho_{c}}$
$=$ $\frac{c\pi}{3}\frac{\text{Im}(\bm{\tau}_{\mu})}{|\bm{\tau}_{\mu}|^{2}\sqrt{1+\frac{4\pi^{2}}{|\bm{\tau}_{\mu}|^{2}}\frac{\rho_{c}}{|\bm{\omega}_{\mu}|^{2}}}},$
which is the same as (4.25).
4.3 Hawking-Page transition and Modular transformation
In AdS3 gravity theory, the Hawking-Page phase transition happens when the free energy of the global AdS equals that of the BTZ black hole and there is only one transition point which is at $|\bm{\tau}|=1$. The free energy of the deformed BTZ and the deformed global AdS are (3.1) and (4.1), which we respectively rewrite here in terms of $(\bm{\omega}_{\mu},\bm{\beta}_{\mu})$:
$I_{E,\text{BTZ}}^{[\mu]}=-\frac{c\pi}{3}\frac{\text{Im}(\bm{\tau}_{\mu})}{|\bm{\tau}_{\mu}|^{2}}\frac{1}{1+\sqrt{1+\frac{4\pi^{2}\rho_{c}}{|\bm{\omega}\mu\bm{\tau}_{\mu}|^{2}}}},$
$I_{E,\text{global}}^{[\mu]}=-\frac{c\pi}{3}\frac{\text{Im}(\bm{\tau}_{\mu})}{1+\sqrt{1+\frac{4\pi^{2}\rho_{c}}{|\bm{\omega}_{\mu}|^{2}}}}.$
By equating them, we find that Hawking-Page phase transition point remains unchanged at $|\bm{\tau}_{\mu}|=1$. Clearly, the deformed on-shell actions $I_{E,\text{global}}^{[\mu]}$ and $I_{E,\text{BTZ}}^{[\mu]}$ are interrelated through the following modular transformation
$\bm{\tau}_{\mu}\leftrightarrow-\frac{1}{\bm{\tau}_{\mu}},\quad\rho_{c}\leftrightarrow\frac{\rho_{c}}{|\bm{\tau}_{\mu}|^{2}}.$ (4.32)
This transformation aligns with the modular property exhibited by the $\overline{\text{T}}$ deformed torus partition function, as proved in [29]. To further support our argument regarding modular transformation, let us proceed to compute the deformed on-shell action for other $SL(2,\mathbb{Z})$ black holes. It is noteworthy that since the $\overline{\text{T}}$ deformation alters the temperature, the modular transformation does not commute with the $\overline{\text{T}}$ deformation. In other words,
$\bm{\tau}_{\mu}(\bm{\tau})\neq-\frac{1}{\bm{\tau}_{\mu}(-\frac{1}{\bm{\tau}})}.$ (4.33)
Therefore, it is not convenient to first construct the $SL(2,\mathbb{Z})$ black holes and then apply the $\overline{\text{T}}$ deformation. Instead, we will directly construct the $\overline{\text{T}}$ deformed $SL(2,\mathbb{Z})$ black holes following the approach outlined in [102]. Starting from the deformed global AdS metric (3.10) with the following two complex periods [^11] Here, we introduce the subscript $m$ to denote the modular transformation and distinguish $c_{m}$ from the CFT central charge $c$.
$\bm{\omega}^{\prime}_{\mu}=\bm{\omega}_{\mu},\quad\bm{\beta}^{\prime}_{\mu}=\bm{\omega}_{\mu}\frac{a_{m}\bm{\tau}_{\mu}+b_{m}}{c_{m}\bm{\tau}_{\mu}+d_{m}},\quad\begin{pmatrix}a_{m}&b_{m}\\ c_{m}&d_{m}\end{pmatrix}\in SL(2,\mathbb{Z}),\quad c_{m}\geq 0,$ (4.34)
and making the following transformation
$V=\frac{1}{c_{m}\bm{\tau}_{\mu}+d_{m}}\hat{V},\quad z=\frac{\hat{z}}{|c_{m}\bm{\tau}_{\mu}+d_{m}|},\quad\rho_{c}=\frac{\hat{\rho_{c}}}{|c_{m}\bm{\tau}_{\mu}+d_{m}|^{2}},$ (4.35)
we end up with the deformed $SL(2,\mathbb{Z})$ black hole metric which is in the form (2) with
$\hat{L}_{0}=-\frac{\pi^{2}}{\bm{\omega}_{\mu}^{2}(c_{m}\bm{\tau}_{\mu}+d_{m})^{2}},\quad\bar{\hat{L}}_{0}=-\frac{\pi^{2}}{\bm{\bar{\omega}}_{\mu}^{2}(c_{m}\bm{\bar{\tau}}_{\mu}+d_{m})^{2}}.$ (4.36)
The two complex periods of the deformed $SL(2,{\mathbb{Z}})$ black hole are given by
$\bm{\hat{\omega}}_{\mu}=\bm{\omega}_{\mu}(a_{m}\bm{\tau}_{\mu}+b_{m}),\quad\bm{\hat{\beta}}_{\mu}=\bm{\omega}_{\mu}(c_{m}\bm{\tau}_{\mu}+d_{m})$ (4.37)
as expected for the $SL(2,{\mathbb{Z}})$ black hole. Since the coordinate transformation (4.35) will not change the on-shell action, we conclude that
$I_{E,SL(2,{Z})}^{[\mu]}=I_{E,\text{global}}^{[\mu]}\left[\frac{a_{m}\bm{\tau}_{\mu}+b_{m}}{c_{m}\bm{\tau}_{\mu}+d_{m}}\Big{|}\frac{\hat{\rho}_{c}}{|c_{m}\bm{\tau}_{\mu}+d_{m}|^{2}}\right]=-\frac{c\pi}{3}\frac{\text{Im}(\frac{a_{m}\bm{\tau}_{\mu}+b_{m}}{c_{m}\bm{\tau}_{\mu}+d_{m}})}{1+\sqrt{1+\frac{4\pi^{2}\hat{\rho}_{c}}{|\bm{\omega}_{\mu}|^{2}|c_{m}\bm{\tau}_{\mu}+d_{m}|^{2}}}},$ (4.38)
which is exactly the modular transformation found in [29]. From the holographic perspective, the modular transformation can be readily understood through the transformation given in equation (4.35). The deformation is irrelevant, and as a result, the deformation parameter $\mu$ has a dimension and scales with the radial coordinate $z$ under a Weyl transformation.
Analogous to the undeformed theory, a modular-invariant semi-classical partition function can be constructed by summing over all bulk saddle solutions [104, 105, 106, 107]:
$Z^{[\mu]}_{\text{torus}}=\sum_{\text{saddles}}e^{-I^{[\mu]}_{E,\text{saddles}}},$ (4.39)
as the modular transformation simply maps one saddle to another. However, this semi-classical partition function is not holomorphic factorized because $\tau_{\mu}$ and $\bar{\tau}_{\mu}$ transform collectively. To further illustrate that the modular property persists even when the holomorphic and anti-holomorphic sectors are asymmetric, we will explore the $\overline{\text{T}}$ deformation of topologically massive gravity, which is dual to a conformal field theory with distinct holomorphic and anti-holomorphic central charges, i.e. $c_{L}\neq c_{R}$[94, 95]. The AdS/CFT correspondence for TMG is elegantly established in [108]. For our purposes of computing the deformed on-shell action, we solely require the expression of the holographic energy-stress tensor to define the $\overline{\text{T}}$ deformation, as we shall demonstrate subsequently.
5 Topologically massive gravity theory
We consider the Euclidean version of the TMG theory whose action is given by
$I_{\text{TMG}}$ $=$ $-\frac{1}{16\pi G_{N}}\int d^{3}x(\mathcal{L}_{EH}+\mathcal{L}_{CS}),$ (5.1)
$\mathcal{L}_{CS}$ $=$ $\frac{\alpha}{2}\sqrt{g}\epsilon^{\lambda\mu\nu}\Gamma^{\rho}_{\lambda\sigma}\left(\partial_{\mu}\Gamma^{\sigma}_{\rho\nu}+\frac{2}{3}\Gamma^{\sigma}_{\mu\tau}\Gamma^{\tau}_{\nu\rho}\right),$ (5.2)
where $\mathcal{L}_{CS}$ is the gravitational Chern-Simons term. Notably, we assume the parameter $\alpha$ to be real, ensuring that the deformed metric remains real [^12] For the Euclidean TMG, derived through a Wick rotation from the Lorentzian theory, the parameter $\alpha$ is purely imaginary because $\mathcal{L}_{CS}$ has odd parity.. The equation of motion for this theory reads
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R-g_{\mu\nu}+{\alpha}C_{\mu\nu}=0,$ (5.3)
where $C_{\mu\nu}$ is the Cotton tensor, defined as
$C_{\mu\nu}=\epsilon_{\mu}^{\alpha\beta}\nabla_{\alpha}(R_{\beta\nu}-\frac{1}{4}g_{\beta\nu}R).$ (5.4)
Consequently, the Einstein solutions (2.11) also satisfy the equations of TMG. However, the holographic stress tensor is modified to [95, 103]
$T_{ij}=\frac{1}{8\pi G_{N}}\left(g_{ij}^{(2)}-g_{ij}^{(0)}g_{kl}^{(2)}g^{kl}_{(0)}\right)+\frac{\alpha}{16\pi G}g^{kl}_{(0)}\left(\epsilon_{lj}g_{ik}^{(2)}+\epsilon_{li}g_{jk}^{(2)}\right),$ (5.5)
where $\epsilon_{ij}$ is the two-dimensional Levi-Civita tensor. The actions of global AdS and BTZ black holes in TMG are given in [109, 111]:
$I_{\text{TMG},\text{global}}^{[0]}=\frac{\text{i}\pi k}{2}\bm{\tau}-\frac{\text{i}\pi\bar{k}}{2}\bm{\bar{\tau}},\quad I_{\text{TMG},\text{BTZ}}^{[0]}=-\frac{\text{i}\pi k}{2\bm{\tau}}+\frac{\text{i}\pi\bar{k}}{2\bm{\bar{\tau}}},$ (5.6)
where the involved parameter $k$ and $\bar{k}$ is determined by the relation
$k-\bar{k}=\frac{\text{i}\alpha}{2G_{N}},\quad k+\bar{k}=\frac{1}{2G_{N}}.$ (5.7)
In the Appendix C, we give a simple derivation [^13] The computation of the on-shell action is a little subtle. For example, one can not obtain the consistent results by substituting the BTZ solution (4.1) into the Euclidean action of the TMG theory. In [109, 111], the on-shell actions were derived in an indirect way. of them based on the Smarr relation.
5.1 $\overline{\text{T}}$ deformation
The $\overline{\text{T}}$ deformation operator is still defined by
$\mathcal{O}_{\text{T$\overline{\text{T}}$ }}=\frac{1}{8}\left(T_{\alpha\beta}T^{\alpha\beta}-(T^{\alpha}_{\alpha})^{2}\right)=T_{vv}T_{\bar{v}\bar{v}}-T_{v\bar{v}}^{2},$ (5.8)
therefore the flow equation is still solved by (2.4) and (2.5). Substituting the stress tensor (5.5) into (2.4) leads to
$\gamma_{\alpha\beta}^{[\mu]}dx^{\alpha}dx^{\beta}=(dv+\frac{\mu\bar{k}}{4\pi}\bar{L}_{0}d\bar{v})(d\bar{v}+\frac{\mu k}{4\pi}L_{0}dv)\equiv dVd\bar{V},$ (5.9)
which implies that
$V=v+\frac{\mu\bar{k}}{4\pi}\bar{L}_{0}\bar{v},\quad\bar{V}=\bar{v}+\frac{\mu k}{4\pi}L_{0}v,$ (5.10)
and
$\bm{\omega}_{\mu}=\bm{\omega}+\frac{\mu\bar{k}}{4\pi}\bar{L}_{0}\bar{\bm{\omega}},\quad\bm{\beta}_{\mu}=\bm{\beta}+\frac{\mu\bar{k}}{4\pi}\bar{L}_{0}\bar{\bm{\beta}}.$ (5.11)
Notice that (5.9) can not be interpreted as an induced metric on a cut-off surface, hence, the cut-off (or glue-on) proposal for $\overline{\text{T}}$ deformation is no longer valid. The deformed on-shell action consists of two parts (2.9) which are convenient to express in terms of quantities in the seed theory:
$I_{\text{TMG}}^{[\mu]}=I_{E}^{[0]}-\mu A_{0}\frac{k\bar{k}}{4\pi^{2}}L_{0}\bar{L}_{0},$ (5.12)
where $A_{0}$ is the area of the seed torus. In the end, we will use the dictionary (5.10) to rewrite it in terms of the quantities in the deformed theory. Until now, our discussion is very general. It is enough to demonstrate the modular transformation between deformed global AdS and deformed BTZ and other case can be generated trivially as we did in section 4.3.
In this case, the global AdS solution corresponds to
$L_{0}=\bar{L}_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}},$ (5.13)
and we find that the deformed on-shell action is
$I_{\text{TMG},\text{global}}^{[\mu]}$ $=$ $\frac{\text{i}\pi}{2}(k\bm{\tau}-\bar{k}\bm{\bar{\tau}})+\frac{\text{i}}{2}\mu(\bm{\tau}-\bm{\bar{\tau}})\frac{|k|^{2}\pi^{2}}{4|\bm{\omega}|^{2}},$
$=$ ${\text{i}\pi}k\frac{\bm{\tau}_{\mu}}{1-\frac{(k-\bar{k})\pi\mu}{4|\bm{\omega}_{\mu}|^{2}}+\sqrt{\frac{k\pi\mu}{|\bm{\omega}\mu|^{2}}+(1-\frac{(k-\bar{k})\pi\mu)}{4|\bm{\omega}_{\mu}|^{2}})^{2}}}$
-- $\text{i}\pi\bar{k}\frac{\bm{\bar{\tau}}_{\mu}}{1+\frac{(k-\bar{k})\pi\mu}{4|\bm{\omega}_{\mu}|^{2}}+\sqrt{\frac{\bar{k}\pi\mu}{|\bm{\omega}_{\mu}|^{2}}+(1+\frac{(k-\bar{k})\pi\mu)}{4|\bm{\omega}_{\mu}|^{2}})^{2}}}.$ (5.15)
Moreover, the BTZ solution corresponds to
$L_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}\bm{\tau}^{2}},\quad\bar{L}_{0}=-\frac{\pi^{2}}{\bm{\bar{\omega}}^{2}\bm{\bar{\tau}}^{2}}.$ (5.16)
Then, we find that the deformed on-shell action is
$I_{\text{TMG},\text{BTZ}}^{[\mu]}$ $=$ $-\frac{\text{i}\pi k}{2\bm{\tau}}+\frac{\text{i}\pi\bar{k}}{2\bm{\bar{\tau}}}+\frac{\text{i}}{2}\mu(\bm{\tau}-\bm{\bar{\tau}})\frac{|k|^{2}\pi^{2}}{4|\bm{\omega}|^{2}|\bm{\tau}|^{4}},$
$=$ $-\frac{\text{i}k\pi}{\bm{\tau}_{\mu}}\frac{1}{1-\frac{(k-\bar{k})\pi\mu}{4|\bm{\omega}_{\mu}|^{2}|\bm{\tau}_{\mu}|^{2}}+\sqrt{\frac{k\pi\mu}{|\bm{\omega}_{\mu}|^{2}|\bm{\tau}_{\mu}|^{2}}+(1-\frac{(k-\bar{k})\pi\mu)}{4|\bm{\omega}_{\mu}|^{2}|\bm{\tau}_{\mu}|^{2}})^{2}}}$
$+$ $\frac{\text{i}\bar{k}\pi}{\bm{\bar{\tau}}_{\mu}}\frac{1}{1+\frac{(k-\bar{k})\pi\mu}{4|\bm{\omega}_{\mu}|^{2}|\bm{\tau}_{\mu}|^{2}}+\sqrt{\frac{\bar{k}\pi\mu}{|\bm{\omega}_{\mu}|^{2}|\bm{\tau}_{\mu}|^{2}}+(1+\frac{(k-\bar{k})\pi\mu)}{4|\bm{\omega}_{\mu}|^{2}|\bm{\tau}_{\mu}|^{2}})^{2}}}.$ (5.18)
Clearly, these two deformed on-shell actions are related by the same modular transformation:
$\bm{\tau}_{\mu}\leftrightarrow-\frac{1}{\bm{\tau}_{\mu}},\quad\mu\leftrightarrow\frac{\mu}{|\bm{\tau}_{\mu}|^{2}}.$ (5.19)
It has been demonstrated in Section 4.3, the change of the deformation parameter under the modular transformation can be understood from the fact that the $\overline{\text{T}}$ deformation is irrelevant. We proceed to investigate the modular properties of the root-$\overline{\text{T}}$ deformation, which is classically marginal. It is worth noting that the root-$\overline{\text{T}}$ deformation is both compatible with the $\overline{\text{T}}$ deformation and also preserves classical integrability [83]. Although the quantum aspects of root-$\overline{\text{T}}$ deformation are less understood [^14] See [84] for some recent progress., a holographic description has been proposed in [85], which bears similarity to the mixed-boundary-condition description of the $\overline{\text{T}}$ deformation. In particular, the spectrum of the deformed theory is accurately reproduced. The authors of [85] also suggest that the root-$\overline{\text{T}}$ deformed CFT might possess certain modular properties. In particular, in [79] it is shown that S-modular transformation is preserved. We provide some evidences by evaluating the deformed on-shell action, as we did for the $\overline{\text{T}}$ deformation.
6 Root-$\overline{\text{T}}$ deformation
Similar to the $\overline{\text{T}}$ deformation, Root-$\overline{\text{T}}$ deformation is also defined by a flow equation
$\frac{dS^{(\mu)}}{d\mu}=\int d^{2}x\sqrt{\gamma}\mathcal{R},\quad\mathcal{R}=\sqrt{\frac{1}{2}T^{\alpha\beta}T_{\alpha\beta}-\frac{1}{4}(T^{\alpha}_{\alpha})^{2}}.$ (6.1)
The operator $\mathcal{R}$ is a multi-trace operator, its holographic dual also corresponds to a deformation of the boundary metric. Because the operator is marginal, it is no need to add a boundary term to the on-shell action like the one in (2.9). The solution of the flow equation is not unique but it is argued in [85], the proper root-$\overline{\text{T}}$ deformed boundary metric and stress tensor are given by
$\gamma_{\alpha\beta}^{(\mu)}=\cosh(\mu)\gamma_{\alpha\beta}^{(0)}+\frac{\sinh(\mu)}{\mathcal{R}^{(0)}}\tilde{T}_{\alpha\beta}^{(0)},\quad\tilde{T}^{(\mu)}_{\alpha\beta}=\cosh(\mu)\tilde{T}_{\alpha\beta}^{(0)}+\sinh(\mu)\mathcal{R}^{(0)}\gamma_{\alpha\beta}^{(0)},$ (6.2)
where $\tilde{T}_{\alpha\beta}$ is the traceless part of the stress tensor $\tilde{T}_{\alpha\beta}=T_{\alpha\beta}-\frac{1}{2}T\gamma_{\alpha\beta}$. We will again start from the seed solution (2.11), for which the stress tensor is $8\pi G_{N}T_{\alpha\beta}=\text{diag}(L_{0},\bar{L}_{0}),\,\tilde{T}_{\alpha\beta}=T_{\alpha\beta}$ and the operator $\mathcal{R}^{(0)}$ equals
$\mathcal{R}^{(0)}=\frac{\sqrt{L_{0}\bar{L}_{0}}}{4\pi G_{N}}.$ (6.3)
Then it is easy to obtain the deformed boundary metric
$\gamma_{\alpha\beta}^{(\mu)}=\cosh(\mu)g_{\alpha\beta}^{(0)}+\frac{\sinh(\mu)}{2\sqrt{L_{0}\bar{L}_{0}}}g_{\alpha\beta}^{(2)},$ (6.4)
and the deformed stress tensor
$\tilde{T}^{(\mu)}_{\alpha\beta}=\frac{1}{8\pi G_{N}}\left(\cosh(\mu)g_{\alpha\beta}^{(2)}+2\sqrt{L_{0}\bar{L}_{0}}\sinh(\mu)g_{\alpha\beta}^{(0)}\right),\quad T^{(\mu)}=T^{(0)},$ (6.5)
where we have used the fact that for a classically marginal deformation, the trace of the stress tensor does not flow. Again we can introduce the flat coordinates on the deformed boundary metric
$\gamma_{\alpha\beta}^{(\mu)}dx^{\alpha}dx^{\beta}=dVd\bar{V}$ (6.6)
where
$dV=\cosh(\frac{\mu}{2})dv+\sqrt{\frac{\bar{L}_{0}}{L_{0}}}\sinh(\frac{\mu}{2})d\bar{v},\quad d\bar{V}=\cosh(\frac{\mu}{2})d\bar{v}+\sqrt{\frac{L_{0}}{\bar{L}_{0}}}\sinh(\frac{\mu}{2})dv,$
$dv=\cosh(\frac{\mu}{2})dV-\sqrt{\frac{\bar{L}_{0}}{L_{0}}}\sinh(\frac{\mu}{2})d\bar{V},\quad d\bar{v}=\cosh(\frac{\mu}{2})d\bar{V}-\sqrt{\frac{L_{0}}{\bar{L}_{0}}}\sinh(\frac{\mu}{2})dV.$
In global AdS3 solution, $L_{0}=\bar{L}_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}}$ then the two periods of the deformed torus are
$\bm{\omega}_{\mu}=\left(\cosh(\frac{\mu}{2})-\sinh(\frac{\mu}{2})\right)\bm{\omega}=e^{-\frac{\mu}{2}}\bm{\omega},\quad\bm{\beta}_{\mu}=\cosh(\frac{\mu}{2})\bm{\beta}+\sinh(\frac{\mu}{2})\bm{\bar{{\beta}}}.$ (6.7)
The deformed on-shell action is given by the first term of (2.16):
$I^{[\mu]}_{E,global}=-\frac{c}{6\pi}\bm{\omega}_{\mu}^{2}\text{Im}(\bm{\tau}_{\mu})\frac{\pi^{2}}{\bm{\omega}_{\mu}^{2}e^{\mu}}=-e^{-\mu}\frac{c\pi}{6}\text{Im}(\bm{\tau}_{\mu}).$ (6.8)
In the BTZ solution with two complex periods $\bm{\omega}=\omega_{1}+\text{i}\omega_{2},\bm{\beta}=\beta_{1}+\text{i}\beta_{2}$, we have
$L_{0}=-\frac{\pi^{2}}{\bm{\beta}^{2}},\quad\bar{L}_{0}=-\frac{\pi^{2}}{\bm{\bar{\beta}^{2}}}.$ (6.9)
Then the deformed periods are
$\bm{\omega}_{\mu}=\cosh\frac{\mu}{2}\bm{\omega}-\sinh\frac{\mu}{2}\frac{\bm{\beta}\bm{\bar{\omega}}}{\bm{\bar{\beta}}},$ (6.10)
$\bm{\beta}_{\mu}=e^{-\frac{\mu}{2}}\bm{\beta}.$ (6.11)
We find that the deformed on-shell action is
$I^{[\mu]}_{\text{E,BTZ}}=-\frac{c}{6\pi}\bm{\omega}_{\mu}^{2}\text{Im}(\bm{\tau}_{\mu})\frac{\pi^{2}}{|\bm{\beta_{\mu}}|^{2}e^{\mu}}=e^{-\mu}\frac{c\pi}{6}\frac{\text{Im}(\bm{\tau}_{\mu})}{|\bm{\tau}_{\mu}|^{2}},$ (6.12)
which is related to $I^{[\mu]}_{\text{E,global}}$ simply by $\bm{\tau}_{\mu}\leftrightarrow-\frac{1}{\bm{\tau}_{\mu}}$. Therefore, our on-shell action calculation suggests that the modular transformation of the root-$\overline{\text{T}}$ deformed CFT is the same as the undeformed one.
7 Discussion
In this work, we study the modular properties of the (root-) $\overline{\text{T}}$ deformed torus partition function of a 2d CFT (with a gravitational anomaly) from the holographic perspective by computing the on-shell actions of various saddle solutions of the dual gravity theories.
For a 2d CFT, we reproduce the modular transformation (2.17). From the bulk point of view, the transformation rule of the deformation parameter $\mu$ manifests due to its dimension being the same as that of the radial coordinate. We also find that the modular property preserves even when the CFT has a gravitational anomaly and the root-$\overline{\text{T}}$ deformation does not alter the modular transformation. Deriving these properties from the field theory side using the same strategy as the one in [29] would be very interesting. Another closely related possible future direction is to study the modular property of the $\text{J}\bar{\text{T}}$ deformed partition function holographically in the mixed-boundary-condition proposal [112]. The $\text{J}\bar{\text{T}}$ deformed partition function is covariant under modular transformation as shown in [113], in particular the deformation parameter also transforms so we expect that it can also be easily understood from the bulk point of view.
Here we only focus on the classical contribution to the partition function. It is crucial but challenging to include the quantum corrections. For pure AdS gravity theory, the quantum correction arises from summing over states corresponding to the Virasoro descendants. In a recent work [35], this 1-loop effect is studied by assuming that the only effect of the $\overline{\text{T}}$ deformation is to alter the temperature, and then the 1-loop contribution can be computed using the usual heat kernel method on a BTZ geometry. However, as shown in [89], the asymptotic symmetry algebra of the deformed theory is not a naive Virasoro algebra, but actually a state-dependent one. This requires further study from both the field theory and gravity theory perspectives.
In the end, we want to emphasize that the boundary term (2.10) in the deformed on-shell action and the holographic dictionary (1) deserve further exploration. For example, as suggested in [55], it may modify the holographic entanglement entropy. In an oncoming work [114], we also find that naively applying the dictionary to the $\overline{\text{T}}$ deformed cone solution or banana solution [115] will lead to wrong results.
Acknowledgments
We thank Miao He, Yan Liu, Yang Lei and Huajia Wang for the valuable discussion. JT is supported by the National Youth Fund No.12105289 and funds from the University of Chinese Academy of Sciences (UCAS), the Kavli Institute for Theoretical Sciences (KITS). FO would like to thank Masahiro Nozaki very much for his support. The work of FO was supported by funds from UCAS, KITS and The Chinese Academy of Sciences (CAS) President’s International Fellowship Initiative (PI FI).
Appendix A Hubbard-Stratonovich method of $\overline{\text{T}}$
In this appendix, we show some details about the derivation about holographic interpretation of $\overline{\text{T}}$ deformation mainly following [77]. Let $\mathcal{O}_{A}$ be a collection of local operators in CFT which are “single trace” in the sense it is dual to some simple bulk field. For example, the energy-momentum tensor $T_{\alpha\beta}$ is treated as a single trace operator because it is dual to the boundary metric $\gamma_{\alpha\beta}$. Then the double trace deformation can be schematically defined as
$S=S_{0}+\frac{N^{2}\mu}{2}\int d^{2}x\sqrt{\gamma}\eta^{AB}\mathcal{O}_{A}\mathcal{O}_{B},$ (A.1)
where $\mu$ is the deformation parameter, $\eta^{AB}$ are some constants for specifying a particular double trace operator and $\gamma_{\mu\nu}$ is the metric of the spacetime where the CFT is defined on. When we derive the holographic dual, we can assume that $\gamma_{\alpha\beta}$ is flat and restore it when it is necessary. The deformed generating function is defined as
$e^{-W^{(\mu)}[J^{(\mu)}]}=\int\mathcal{D}\psi\exp\left(-S_{0}-N^{2}\int d^{2}x\left(\frac{\mu}{2}\eta^{AB}\mathcal{O}_{A}\mathcal{O}_{B}+J^{(\mu)A}\mathcal{O}_{A}\right)\right).$ (A.2)
Following Hubbard-Stratonovich, we integrate in the fields $h_{A}$ by inserting the resolution of identity
$1=\mathcal{N}\int\mathcal{D}h\exp\left(\frac{N^{2}}{2\mu}\int d^{2}h^{A}(\eta^{-1})_{AB}h^{B}\right)$ (A.3)
and combine the two quadratic terms which leads to
$e^{-W^{(\mu)}[J^{(\mu)}]}\!\!\!$ $\!\!\!\!\!=\mathcal{N}\int\mathcal{D}h\mathcal{D}\psi\exp\left(-S_{0}+N^{2}\int d^{2}x\frac{1}{2\mu}(h^{A}+\mu\eta^{AC}\mathcal{O}_{C})(\eta^{-1})_{AB}(h^{B}+\mu\eta^{BD}\mathcal{O}_{D})\right.$ (A.4)
$\left.-N^{2}\int d^{2}x\mathcal{O}_{A}(J^{(\mu)A}+\mu\eta^{AB}\mathcal{O}_{B}+h^{A})\right).$
Introducing the shifted source
$\tilde{J}^{A}=J^{(\mu)A}+\mu\eta^{AB}\mathcal{O}_{B}+h^{A},$ (A.5)
then the generating function can be cast into
$e^{-W^{(\mu)}[J^{(\mu)}]}=\mathcal{N}\int\mathcal{D}h\,e^{-W^{(0)}[\tilde{J}]}\exp\left(\frac{N^{2}}{2\mu}\int d^{2}x(\tilde{J}^{A}-J^{(\mu)A})(\eta^{-1})_{AB}(\tilde{J}^{B}-J^{(\mu)B})\right).\;\;\;\;\;\;\;\;$ (A.6)
Ultimately, we evaluate the $h$ integral with saddle point approximation. Varying the exponent with respect to $h_{A}$ gives
$-\mathcal{O}_{A}+\frac{1}{\mu}(\eta^{-1})_{AB}(\tilde{J}^{B}-J^{(\mu)B})=0,$ (A.7)
with the solution
$\tilde{J}^{A}=\mu\eta^{AB}\mathcal{O}_{B}+J^{(\mu)A},$ (A.8)
Substituting it into (A.6) we end up with
$W^{(\mu)}[J^{(\mu)}]=W^{(0)}[J^{(\mu)A}+\mu\eta^{AB}\mathcal{O}_{B}]-\frac{N^{2}\mu}{2}\int d^{2}x\eta_{AB}\mathcal{O}_{A}\mathcal{O}_{B}.$ (A.9)
Note that this is only the result of the first-order deformation. To obtain the deformed theory on the whole $\overline{\text{T}}$ flow we have to work out 1) the deformed source $J^{(\mu)}$; 2) the deformed functional $W^{(\mu)}$. The deformed source $J^{(\mu)}$ has been derived by the variation method we review below. Using the defining property of the generating function $\delta W=\frac{1}{2}\int d^{2}x\sqrt{\gamma}T_{\alpha\beta}\delta\gamma^{\alpha\beta}$, we can obtain a flow equation
$\frac{1}{2}\partial_{\mu}\left(\int d^{2}x\sqrt{\gamma^{[\mu]}}T_{\alpha\beta}^{[\mu]}\delta\gamma_{\alpha\beta}^{[\mu]}\right)=-\delta\left(\int d^{2}x\sqrt{\gamma^{[\mu]}}\text{T}\bar{\text{T}}^{[\mu]}\right),$ (A.10)
which can be solved by
$\gamma_{\alpha\beta}^{[\mu]}=\gamma_{\alpha\beta}^{[0]}+\frac{1}{2}\mu\hat{T}_{\alpha\beta}^{[0]}+\frac{1}{16}\mu^{2}\hat{T}_{\alpha\rho}^{[0]}\hat{T}_{\sigma\beta}^{[0]}{\gamma^{[0]}}^{\rho\sigma},$ (A.11)
$\hat{T}^{[\mu]}_{\alpha\beta}=\hat{T}^{[0]}_{\alpha\beta}+\frac{1}{4}\mu\hat{T}_{\alpha\rho}^{[0]}\hat{T}_{\sigma\beta}^{[0]}{\gamma^{[0]}}^{\rho\sigma},$ (A.12)
where $\hat{T}_{\alpha\beta}=T_{\alpha\beta}-\gamma_{\alpha\beta}T$. Some details are the following. Explicitly, the flow equation is
$\partial_{\mu}(\sqrt{\gamma}T_{\alpha\beta}\delta\gamma^{\alpha\beta})=-\frac{1}{4}\delta(\sqrt{\gamma}(T_{\alpha\beta}T^{\alpha\beta}-(T^{\alpha}_{\alpha})^{2}))\equiv\delta(\sqrt{\gamma}\mathcal{O}_{\text{T$\overline{\text{T}}$ }}).$ (A.13)
The left-hand-side is
$\partial_{\mu}(\sqrt{\gamma}T_{\alpha\beta})\delta\gamma^{\alpha\beta}+\sqrt{\gamma}T_{\alpha\beta}\delta(\partial_{\mu}\gamma^{\alpha\beta}),$ (A.14)
and the right-hand-side is
$-\frac{1}{4}\sqrt{\gamma}\left(2(T_{\alpha\beta}\delta T^{\alpha\beta}+T^{\alpha\gamma}T^{\beta}_{\gamma}\delta\gamma_{\alpha\beta}-T\delta T)\right)-\frac{1}{2}\sqrt{\gamma}\mathcal{O}_{\text{T$\overline{\text{T}}$ }}\gamma_{\alpha\beta}\delta\gamma^{\alpha\beta}$ (A.15)
$=\sqrt{\gamma}\left(C_{1}\delta\gamma^{\alpha\beta}+T_{\alpha\beta}\delta C_{2}\right).$ (A.16)
Comparing the lhs and rhs, we get
$\partial_{\mu}(\sqrt{\gamma}T_{\alpha\beta})=C_{1}\sqrt{\gamma},$ (A.17)
$\partial_{\mu}\gamma^{\alpha\beta}=C_{2},$ (A.18)
with
$C_{1}=-\frac{1}{2}\mathcal{O}_{\text{T$\overline{\text{T}}$ }}\gamma_{\alpha\beta}+\frac{1}{2}T_{\alpha}^{\delta}T_{\delta\beta}-\frac{1}{2}TT_{\alpha\beta},$ (A.19)
$C_{2}=-\frac{1}{2}T^{\alpha\beta}+\frac{1}{2}T\gamma^{\alpha\beta}.$ (A.20)
The proposal is that $\gamma^{[\mu]}_{\alpha\beta}$ and $T^{[\mu]}_{\alpha\beta}$ are the sources and dual operators of the deformed holographic theory which is still an asymptotic AdS3 gravity theory. The general asymptotic AdS3 solution can be written in the Fefferman-Graham gauge like
$ds^{2}=g_{\alpha\beta}(\rho,x^{\alpha})dx^{\alpha}dx^{\beta}+\frac{d\rho^{2}}{4\rho^{2}},\quad g_{\alpha\beta}(\rho,x^{\alpha})=\frac{g_{\alpha\beta}^{(0)}}{\rho}+g_{\alpha\beta}^{(2)}+\rho g_{\alpha\beta}^{(4)},$ (A.21)
where $g^{(2)}$ corresponds to the initial expectation value of the CFT operator
$\hat{T}^{[0]}_{\alpha\beta}=\frac{1}{8\pi G_{N}}g_{\alpha\beta}^{(2)}.$ (A.22)
The solution (A.11) implies that the deformed boundary metric is
$\gamma_{\alpha\beta}^{[\mu]}=g_{\alpha\beta}^{(0)}+\frac{\mu}{16\pi G_{N}}g_{\alpha\beta}^{(2)}+\frac{\mu^{2}}{(16\pi G_{N})^{2}}g_{\alpha\beta}^{(4)}=\rho_{c}g_{\alpha\beta}(\rho_{c}),\quad\rho_{c}=\frac{\mu}{16\pi G_{N}}.$ (A.23)
To derive the deformed functional, we notice
$W^{(\mu)}$ $=$ $W^{(\mu-\delta\mu)}[J^{(\mu-\delta\mu)}+\delta\mu\eta^{AB}\mathcal{O}^{(\mu-\delta\mu)}]-\delta\mu\frac{N^{2}}{2}\int d^{2}x\eta_{AB}[\mathcal{O}_{A}\mathcal{O}_{B}]^{(\mu)}$ (A.24)
[…]
In the case of $\overline{\text{T}}$ we have
[…]
$=$ $W^{(0)}[\gamma_{\alpha\beta}^{[\mu]}=g_{\alpha\beta}^{(0)}+\frac{\mu}{16\pi G_{N}}g_{\alpha\beta}^{(2)}+\frac{\mu^{2}}{(16\pi G_{N})^{2}}g_{\alpha\beta}^{(4)}]-\mu\int d^{2}x\sqrt{\gamma^{(0)}}\text{T$\overline{\text{T}}$ }^{(0)}.$
In the last step we have used the fact $\sqrt{\gamma^{(k)}}\text{T$\overline{\text{T}}$ }^{(k)}$ is a flow invariant and
$\sqrt{\gamma^{(\mu)}}\text{T$\overline{\text{T}}$ }^{(\mu)}=\sqrt{\gamma^{(0)}}\text{T$\overline{\text{T}}$ }^{(0)}.$ (A.27)
Appendix B BTZ black holes with complex periods
It is straightforward to transform the metric (4.1) into the FG gauge (2.11) with
$L_{0}$ $=$ $\frac{\left(R_{+}+\text{i}R_{-}\right)^{2}}{4}=\frac{\pi^{2}}{\left(\beta-\text{i}\theta\right)^{2}}\equiv-\frac{\pi^{2}}{\bm{\beta}^{2}},$ (B.1)
$\bar{L}_{0}$ $=$ $\frac{\left(R_{+}-\text{i}R_{-}\right)^{2}}{4}=\frac{\pi^{2}}{\left(\beta+\text{i}\theta\right)^{2}}\equiv-\frac{\pi^{2}}{\bar{\bm{\beta}}^{2}}.$
From the holographic stress tensor (A.22), one can derive the holographic mass and momentum
[…]
The entropy of the black hole is equal to
$S_{BH}=\frac{2\pi R_{+}}{4G_{N}}=\frac{\pi c}{3}\left(\sqrt{L_{0}}+\sqrt{\bar{L}_{0}}\right)=\frac{\pi^{2}c}{3}\frac{2\beta}{\beta^{2}+\theta^{2}}.$ (B.4)
Now to obtain a generic torus, we do a coordinate transformation
$v\rightarrow\frac{\omega}{2\pi}e^{\text{i}\alpha}v=\hat{v},$ (B.5)
or equivalently
$\begin{pmatrix}\hat{\phi}\\ \hat{t}\end{pmatrix}=\frac{\omega}{2\pi}\begin{pmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{pmatrix}\begin{pmatrix}\phi\\ t\end{pmatrix},$ (B.6)
which implies that the periods are transformed as
$\begin{pmatrix}\hat{\omega}_{1}\\ \hat{\omega}_{2}\end{pmatrix}$ $=$ $\frac{\omega}{2\pi}\begin{pmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{pmatrix}\begin{pmatrix}\omega_{1}\\ \omega_{2}\end{pmatrix},$ (B.7)
$\begin{pmatrix}\hat{\beta}_{1}\\ \hat{\beta}_{2}\end{pmatrix}$ $=$ $\frac{\omega}{2\pi}\begin{pmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{pmatrix}\begin{pmatrix}\beta_{1}\\ \beta_{2}\end{pmatrix},$ (B.8)
where $\omega_{1}=2\pi,\omega_{2}=0,\omega_{1}+\text{i}\omega_{2}=\bm{\omega},$ and $\beta_{1}=\theta,\beta_{2}=\beta,\beta_{1}+\text{i}\beta_{2}=\bm{\beta}$. Noting the this transformation is a combination of Weyl transformation and rotation, the holographic mass and momentum should transform accordingly as
$\begin{pmatrix}\hat{J}\\ \hat{M}\end{pmatrix}=\frac{2\pi}{\omega}\begin{pmatrix}J\cos\alpha-M\sin\alpha\\ J\sin\alpha+M\cos\alpha\end{pmatrix},$ (B.9)
which lead to (4.5) and (4.6). In addition, these two quantities correspond to the conserved charges of $\partial_{\hat{\phi}}$ and $\partial_{\hat{t}}$ respectively.
Appendix C On-Shell Action of TMG Theory
In this appendix, we will derive the on-shell action (5.6) of TMG theory by assuming the Smarr relation still holds. The Smarr relation can be viewed as a relationship between energy and the entropy and it holds separately for the holomorphic and anti-holomorphic sectors. Therefore, it is not modified even when $c_{L}\neq c_{R}$. Starting from the following expressions
$T_{vv}=\frac{1+\text{i}\alpha}{8\pi G_{N}}L_{0}=\frac{k}{2\pi}L_{0},\quad T_{\bar{v}\bar{v}}=\frac{1-\text{i}\alpha}{8\pi G_{N}}\bar{L}_{0}=\frac{\bar{k}}{2\pi}\bar{L}_{0},$ (C.1)
it is straightforward to obtain the holographic mass and momentum
[…]
here we have chosen a real $\omega$. Actually, with the help of (4.5) and (4.6), it is straightforward to obtain the result with complex period $\bm{\omega}$ as follows
$M=\frac{\bm{\omega}kL_{0}+\bar{\bm{\omega}}\bar{k}\bar{L}_{0}}{2\pi},\quad J=\frac{-\text{i}(\bm{\omega}kL_{0}-\bar{\bm{\omega}}\bar{k}\bar{L}_{0})}{2\pi}.$ (C.4)
Therefore, for global AdS the on-shell action is given by
$I^{[0]}_{\text{TMG,global}}=\beta_{2}M+\beta_{1}J=\frac{-\text{i}(\bm{\beta}\bm{\omega}kL_{0}-\bar{\bm{\beta}}\bar{\bm{\omega}}\bar{k}\bar{L}_{0})}{2\pi}.$ (C.5)
Next by applying $L_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}}$ and $\bm{\beta}=\bm{\omega}\bm{\tau}$, the above result can be simplified to
$I^{[0]}_{\text{TMG,global}}=\frac{\text{i}\pi k}{2}\bm{\tau}-\frac{\text{i}\pi\bar{k}}{2}\bar{\bm{\tau}}.$ (C.6)
On the other hand, for the BTZ black hole, the on-shell action is as follows
$I^{[0]}_{\text{TMG,BTZ}}=\beta_{2}M+\beta_{1}J-S_{BH}.$ (C.7)
Using the Smarr relation
$2(\beta_{2}M+\beta_{1}J)=S_{BH},$ (C.8)
we find that the on-shell action equals to
$I^{[0]}_{\text{TMG,BTZ}}=-(\beta_{2}M+\beta_{1}J)=-\frac{\text{i}\pi k}{2\bm{\tau}}+\frac{\text{i}\pi\bar{k}}{2\bar{\bm{\tau}}},$ (C.9)
where we have used the fact $L_{0}=-\frac{\pi^{2}}{\bm{\beta}^{2}}$ and $\bar{L}_{0}=-\frac{\pi^{2}}{\bar{\bm{\beta}}^{2}}$.